# Invariants for $\mathbb{G}_{(r)}$-modules

### v1

@@ 4. Exponential Jordan type
By Proposition 2.13, $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ is a coherent, locally free sheaf of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}$-modules on $\mathbb{P}\mathcal{C}_{r}$. Proposition 4.5(3) implies that $(\Phi_{r|s})^{*}:k[V_{r}(\mathbb{G})]\to k[\mathcal{C}_{r}(\mathfrak{g})]$ is a map of graded $k$-algebras so that $\Theta_{\mathfrak{g},r}^{exp}$ is homogenous by Theorem 2.11(5) and Theorem 4.8(1). Thus, $\Theta_{\mathfrak{g},r}^{exp}$ determines $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}$ in
$\Gamma(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})$.
@@
The following is our simplification of Theorem 2.16 for $(\mathbb{G},\mathcal{E}_{r})$ $(\mathbb{G},\mathcal{E}_{r})$ is an affine group scheme of exponential type of height $r$. The proof of Theorem 4.16 follows from Proposition 4.15 exactly as the proof of Theorem 2.16 follows from Theorem 2.14.
@@
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. For any finite dimensional $\mathbb{G}_{(r)}$-module $M$ and any scheme-theoretic point $\underline{B}$ of $\mathcal{C}_{r}(\mathfrak{g})$ projecting to $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$, we define
$JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)\quad\equiv\quad JT(\sum_{s=0}^{p-1}(\mathcal{E}_{(r),B_{s}})_{*}(u_{r-s-1}),k(\underline{B})\otimes M),$
the Jordan type of the $p$-nilpotent element  $k(\underline{B})\otimes_{\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}}\tilde{\Theta}_{\mathfrak{g},r}^{exp}\in k(\underline{B})\otimes\mathbb{G}_{(r)}$ acting on $k(\underline{B})\otimes M$.
@@
    For any $\mathbb{G}_{(r)}$-module $M$ of dimension $m$, the association $\langle\underline{B}\rangle\ \mapsto JT_{\mathfrak{g},r,M}^{exp}(\underline{B})$ determines a continuous map
    $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\quad\to\quad\mathcal{Y}_{\leq,m}.$
@@
    For $\mathbb{G}_{(r)}$-modules $M,N$ of dimensions $m,n$ and any scheme-theoretic point $\underline{B}\in\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$,
    $JT_{\mathfrak{g},r,M\oplus N}^{exp}(\langle\underline{B}\rangle)\ =\ JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)+JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \in\ \mathcal{Y}_{\leq,m+n}.$
@@
    Assume either that $r=1$ or that $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$ is a generic point. For $\mathbb{G}_{(r)}$-modules $M,N$ of dimensions $m,n$,
    $JT_{\mathfrak{g},r,M\otimes N}^{exp}(\langle\underline{B}\rangle)\ =\ JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)\otimes JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \in\ \mathcal{Y}_{\leq,mn}.$
@@
    If $f:M\to N$ is a surjective map of $\mathbb{G}_{(r)}$-modules, then $JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \leq JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)$ for any $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$; in other words, $JT_{\mathfrak{g},r,N}^{exp}(-)\leq JT_{\mathfrak{g},r,M}^{exp}(-)$.
@@
    If $\phi:\mathbb{H}\to\mathbb{G}$ is a morphism of affine group schemes of exponential type of height $r$, then for any $\mathbb{G}_{(r)}$-module $M$ of dimension $m$
    $JT_{\mathfrak{h},r,\phi*(M)}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{h})\ \to\ \mathcal{Y}_{\leq,m}$
@@
    $JT_{\mathfrak{g},r,M}^{exp}(-)\circ\phi_{*}:\mathbb{P}\mathcal{C}_{r}(\mathfrak{h})\ \to\ \mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\ \to\ \mathcal{Y}_{\leq,m}.$
@@ 5. Examples of exponential Jordan types
We provide some computations of exponential Jordan types using the simplified Jordan type function $JT^{exp}_{\mathfrak{g},r,M}(-)$ of Theorem 4.16. One challenge to computing Jordan types is the behavior of Jordan types with respect to the sum of two commuting, $p$-nilpotent operators. Another challenge is the fact that for $r>1$ the Jordan type functions $JT_{\mathbb{G},r,M}(-),\ JT_{\mathfrak{g},r,M}^{exp}(-)$ is unlikely to commute with tensor products of $\mathbb{G}_{(r)}$-modules at all points of $\mathbb{P}V_{r}(\mathbb{G}),\ \mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$.
@@
Example 3.5 remains basically unchanged, since the Jordan type of any $p$-nilpotent $\alpha\in\mathfrak{g}_{K}$ is the Jordan type of $exp(\alpha)$.
On the other hand, as we see in the following example, computations much simplify if we replace $JT_{\mathbb{G}_{(r)},M}(-)$ by $JT^{exp}_{\mathfrak{g},M}(-)$.
@@
Adopt the notation of Examples 2.6 and 2.10. Let $V$ be an $n$-dimensional $k$-vector space, and let $\alpha_{0},\ldots\alpha_{r-1}$ be be pair-wise commuting $p$-nilpotent endomorphisms of $V$. Denote by $M$ the $(\mathbb{G}_{a}^{r})_{1}$-module such that $x_{i}\in k(\mathbb{G}_{a}^{r})_{1}$ acts as $\alpha_{i}$ and denote by $N$ the $\mathbb{G}_{a(r)}$-module such that $u_{i}\in k\mathbb{G}_{a(r)}$ also acts as $\alpha_{i}$.
Let $\overline{x}$ be a scheme-theoretic point of $\mathbb{P}^{r-1}$, let $\overline{y}=\Psi(\overline{x})$, and let $\overline{x}^{(r-1)}=\Phi((\overline{y})=F^{r-1}(\overline{x})$, where $\Psi:w\mathbb{P}(1,p,\ldots,p^{r-1})\to\mathbb{P}^{r-1},\ \Phi:\mathbb{P}^{r-1}\to w\mathbb{P}(1,p,\ldots,p^{r-1})$ are discussed in Example 2.6. Then
$JT_{\mathfrak{g}_{a},r,N}^{exp}(\overline{y})\quad=\quad JT_{\mathbb{G}_{a}^{\times r},1,M}(\overline{x}^{(r-1)}).$
This is verified by observing that we may identify $\Phi^{*}(\Theta_{\mathbb{G}_{a},r}^{exp})$ with $F^{r-1*}(\Theta_{\mathbb{G}_{a},1})$ in $k[V_{1}(\mathbb{G}_{a}^{\times r})]\otimes k[z_{0},\ldots,z_{r-1}]/(z_{i}^{p})$.
Let $(\mathbb{G},\mathcal{E}_{r})$ be associated with an embedding $i:\mathbb{G}\hookrightarrow GL_{N}$ of exponential type of height $r$, so that $\mathcal{E}_{r}:\mathbb{G}_{a(r)}\to\mathbb{G}_{(r)}$ is the restriction of the exponential map for $GL_{N}$. Thus, $JT^{exp}_{\mathfrak{g},r,M}(-):\mathbb{P}\mathcal{C}_{r}(\mathbb{G})\quad\to\quad\mathcal{Y}_{\leq m}$ sends a scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathfrak{g})$ to the Jordan type of $\sum_{s=0}^{r-1}(exp_{B_{s}})_{*}(u_{r-s-1})\in GL_{m}(k(\underline{B}))$ acting on $k(\underline{B})\otimes M$.
@@
Let $\mathbb{G}=SL_{2(2)}$. Consider the irreducible $SL_{2(2)}$-modules $S_{\lambda}$ for $10\leq\lambda\leq p^{2}-$ 1, so that $S_{\lambda}=S(\lambda_{0})\otimes S(\lambda_{1})^{(1)}$ where $0\leq\lambda_{1},\lambda_{2}<p$ and $\lambda=\lambda_{0}+p\lambda_{1}$. What is implicit in this computation is that $u_{i}$ vanishes on $S(\lambda_{0})$ and on $S(\lambda_{1})$ for all $i\geq 1$. The dimensions of $S(\lambda_{0}),S(\lambda_{1})^{(1)},S(\lambda_{0})\otimes S(\lambda_{1})^{(1)}$ are respectively $m=\lambda_{0}+1,n=\lambda_{1}+1$, and $m\cdot n=(\lambda_{0}+1)\cdot(\lambda_{1}+1)$. To compute $JT_{SL_{2},2,S_{\lambda}}(B_{0},B_{1})$ for $B_{0},B_{1}$ commuting elements in $sl_{2}$, we appeal to Proposition 2.18 to reduce to the case that $B_{0}=a_{0}E,\ B_{1}=a_{1}E$ where $E$ is the $2\times 2$ matrix with $E_{1,2}=1$ and $E_{1,1}=E_{2,1}=E_{2,2}=0$.
Thus, to compute $JT_{sl_{2},2,S(\lambda)}^{exp}(-)$ it suffices to compute the Jordan types of $a_{1}^{p}E\cdot u_{0}+a_{0}E\cdot u_{1}$ on $S(\lambda_{0})\otimes S(\lambda_{1})^{(1)}$. This equals the Jordan type of $a_{0}E$ on $S(\lambda_{1})$ tensor (as Jordan types – see Definition 1.5) the Jordan type of $a_{1}^{p}E$ on $S(\lambda_{0})$. Observe that the Jordan type of $E$ on $S(\lambda_{0})=[m]$ and on $S(\lambda_{1})=[n]$. Thus, if both $a_{0},\ a_{1}$ are non-zero, then the the Jordan type of $a_{1}^{p}E\cdot u_{0}+a_{0}E\cdot u_{1}$ on $S(\lambda)$ is $[m]\otimes[n]$; if $a_{0}\not=0,\ a_{1}=0$ then the Jordan type $S(\lambda)$ is $m\cdot[n]$; and if $a_{0}=0,\ a_{1}\not=0$ then the Jordan type $S(\lambda)$ is $n\cdot[m]$
@@
Let $V$ be a polynomial representation of $GL_{N}$ of degree $<p$ and consider the $i$-th Frobenius twist $V^{(i)}$ for some $i\geq 0$. We fix some $r>i$ and proceed to determine $JT_{\mathfrak{g}l_{N},r,V^{(i)}})(-)$. Then the action of $(\mathcal{E}_{B_{s}})_{*}(u_{r-s-1})$ on $V^{(i)}$ is non-trivial if and only if $10\leq s=r-i-$ 1. If $s=r-i-1$ so that $i=r-s-1$, then this action is given by the action of the exponential of $B_{s}$ and thus has Jordan type that of $B_{s}\in\mathfrak{g}l_{N}$ acting on $V$. Consequently, $JT_{\mathfrak{g}l_{N},r,V^{(i)}}^{exp}(\underline{B})=JT(B_{s},V)$ where $s=r-i-1$.
More generally, we can take a polynomial representation $V_{i}$ of $GL_{N}$ for each $i,0\leq i<r$ with each $V_{i}$ of degree $<p$ and consider $M=\otimes_{i=1}^{r-1}V_{i}^{(i)}$. Then we may use the preceding paragraph together with the rule for tensor product of Jordan types to determine $JT_{\mathfrak{g}l_{N},r,M}^{exp}(-)$ as in Example 5.3 in terms of the Jordan types of each $JT(B_{r-i-1},V_{i})$.
@@
As a special case of Example 5.4, we take $V=\mathfrak{g}^{ad}$ to be the adjoint representation (i.e., equal to the Lie algebra $\mathfrak{g}L_{N}$ with conjugate action). In this case, $JT_{\mathfrak{g}l_{N},r,\mathfrak{g}^{ad}}^{exp}(\underline{B})$ is the Jordan type of $[-,B_{0}]:\mathfrak{g}l_{N}\to\mathfrak{g}l_{N}$.
This generalizes to $\mathbb{G}$ with $V=\mathfrak{g}^{ad}$ provided that $i:\mathbb{G}\hookrightarrow GL_{N}$ is an embedding of exponential type of height $r$.
6. Vector bundles
In this section we revisit the construction of vector bundles associated to certain $\mathbb{G}$-modules given in [14]. Although the discussion below applies more generally to $\mathbb{G}$-modules of constant $j$-type, we restrict our discussion to modules of constant Jordan type. We refer to [5] for an introduction to modules of constant Jordan type and for numerous examples of these interesting $\mathbb{G}$-modules.
The Jordan type functions we have considered are “pointwise” in the sense that their value at a point $\underline{B}\in\mathcal{C}_{r}(\mathfrak{g})$ depends upon the specialization of $k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M$ along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathcal{C}_{r}(\mathfrak{g})$. In [14], vector bundles were associated to $\mathbb{G}$-modules of constant Jordan type in part because their invariants can sometimes distinguish two $\mathbb{G}$-modules whose Jordan type functions are equal.
We proceed to verify that these vector bundles (i.e., locally free, coherent sheaves) are little changed if we replace the universal $p$-nilpotent operator $\Theta_{\mathbb{G},r}$ by the simpler $\Theta_{\mathfrak{g},r}^{exp}$.

### v2

@@ 4. Exponential Jordan type
By Proposition 2.17, $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ is a coherent, locally free sheaf of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$. Proposition 4.5(3) implies that $(\Phi_{r|s})^{*}:k[V_{r}(\mathbb{G})]\to k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$ is a map of graded $k$-algebras so that $\Theta_{\mathfrak{g},r}^{exp}$ is homogenous by Theorem 2.15(5) and Theorem 4.8(1). Thus, $\Theta_{\mathfrak{g},r}^{exp}$ determines $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}$ in
$\Gamma(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})$.
@@
[…]
@@
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. For any finite dimensional $\mathbb{G}_{(r)}$-module $M$ and any scheme-theoretic point $\underline{B}$ of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ projecting to $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$, we define
$JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)\quad\equiv\quad JT(\sum_{s=0}^{p-1}(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1}),k(\underline{B})\otimes M),$
the Jordan type of the $p$-nilpotent element $k(\underline{B})\otimes_{\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}}\tilde{\Theta}_{\mathfrak{g},r}^{exp}\in k(\underline{B})\otimes\mathbb{G}_{(r)}$ acting on $k(\underline{B})\otimes M$.
@@
    For any finite dimensional $\mathbb{G}_{(r)}$-module $M$, the association $\langle\underline{B}\rangle\ \mapsto JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)$ determines a continuous map
    $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\quad\to\quad\mathcal{Y}.$
@@
    For $\mathbb{G}_{(r)}$-modules $M,N$ of dimensions $m,n$ and any scheme-theoretic point $\underline{B}\in\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$,
    $JT_{\mathfrak{g},r,M\oplus N}^{exp}(\langle\underline{B}\rangle)\ =\ JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)+JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \in\ \mathcal{Y}_{m+n}.$
@@
    Assume either that $r=1$ or that $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ is a generic point. For $\mathbb{G}_{(r)}$-modules $M,N$ of dimensions $m,n$,
    $JT_{\mathfrak{g},r,M\otimes N}^{exp}(\langle\underline{B}\rangle)\ =\ JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)\otimes JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \in\ \mathcal{Y}_{m\cdot n}.$
@@
    If $f:M\to N$ is a surjective map of $\mathbb{G}_{(r)}$-modules, then $JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \leq JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)$ for any $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$; in other words, $JT_{\mathfrak{g},r,N}^{exp}(-)\leq JT_{\mathfrak{g},r,M}^{exp}(-)$.
@@
    If $\phi:\mathbb{H}\to\mathbb{G}$ is a morphism of affine group schemes of exponential type of height $r$, then for any finite dimensional $\mathbb{G}_{(r)}$-module $M$
    $JT_{\mathfrak{h},r,\phi*(M)}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{h}))\ \to\ \mathcal{Y}$
@@
    $JT_{\mathfrak{g},r,M}^{exp}(-)\circ\phi_{*}:\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{h}\ \to\ \mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ \mathcal{Y}.$
@@ 5. Examples of exponential Jordan types
We provide some examples of computations of the simplified Jordan type function $JT^{exp}_{\mathfrak{g},r,M}(-)$ of Theorem 4.18. One challenge to computing Jordan types is the behavior of Jordan types with respect to the sum of two commuting, $p$-nilpotent operators. Another challenge is the fact that for $r>1$ the Jordan type functions $JT_{\mathbb{G},r,M}(-),\ JT_{\mathfrak{g},r,M}^{exp}(-)$ rarely commute with tensor products of $\mathbb{G}_{(r)}$-modules at all points of $\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$.
@@
Example 3.6 remains unchanged, since $JT_{\mathbb{G},1,M}(-)\ =\ JT^{exp}_{\mathfrak{g},1,M}(-)$.
On the other hand, as we see in the following examples, some computations simplify if we replace $JT_{\mathbb{G}_{(r)},M}(-)$ by $JT^{exp}_{\mathfrak{g},M}(-)$.
@@
Adopt the notation of Examples 2.10 and 2.14. Let $V$ be an $n$-dimensional $k$-vector space, and let $\alpha_{0},\ldots\alpha_{r-1}$ be be pair-wise commuting $p$-nilpotent endomorphisms of $V$. Denote by $M$ the $(\mathbb{G}_{a}^{\times r})_{1}$-module such that $x_{i}\in k(\mathbb{G}_{a}^{r})_{1}$ acts as $\alpha_{i}$ and denote by $N$ the $\mathbb{G}_{a(r)}$-module such that $u_{i}\in k\mathbb{G}_{a(r)}$ also acts as $\alpha_{i}$.
Recall from Example 2.10 the morphism $\Phi:\mathbb{P}^{r-1}\to w\mathbb{P}(1,p,\ldots,p^{r-1})$. Since $\Theta_{\mathbb{G}_{a,r}}^{exp}\ =\ \sum_{i=0}^{r-1}T_{i}^{p^{r-1-i}}\otimes u_{i}\ \in\ k[V_{r}(\mathbb{G}_{a(r)})]\otimes k\mathbb{G}_{a(r)}$,
$\Phi^{*}(\Theta_{\mathbb{G}_{a},r}^{exp})\quad=\quad\sum_{i=0}^{r-1}t_{i}^{p^{r-1}}\otimes x_{i}\ =\ F^{r-1*}(\Theta_{\mathbb{G}_{a}^{\times r},1}).$
Thus,
$JT_{\mathfrak{g}_{a},r,N}^{exp}(-)\circ\Phi\ =\ JT_{\mathbb{G}_{a}^{\times r},1,M}\circ F^{r-1}:\ \mathbb{P}^{r-1}\ \to\ \mathcal{Y}.$
In the following examples, we use (4.16.1) of Definition 4.16 which becomes
$JT^{exp}_{\mathfrak{g},r,M}(\underline{B})\quad=\quad JT(\sum_{s=0}^{r-1}(exp_{B_{s}})_{*}(1\otimes u_{r-s-1}),k(\underline{B}))$
provided that $(\mathbb{G},\mathcal{E}_{r})$ is associated with an embedding $i:\mathbb{G}\hookrightarrow GL_{N}$ of exponential type of height $r$.
@@
Set $\mathbb{G}=SL_{2(2)}$, so that $\mathfrak{g}=sl_{2}$. We consider the irreducible $\mathbb{G}$-module
$M\ =\ S(\lambda_{0})\otimes S(\lambda_{1})^{(1)},\quad 0\leq\lambda_{0},\lambda_{1}<p,\ \lambda=\lambda_{0}+p\lambda_{1}.$
We proceed to compute $JT^{exp}_{sl_{2},2,M}(-):\mathcal{C}_{2}(\mathcal{N}_{p}(sl_{2}))\to\mathcal{Y}_{\leq}$ as in Definition 4.16. To compute $JT_{SL_{2},2,M}(\underline{B})$ for $B_{0},B_{1}$ commuting elements in $sl_{2}$, we appeal to Proposition 2.20 to reduce to the case that $\underline{B}=(a_{0}E,a_{1}E)$ where $E$ is the $2\times 2$ matrix with $E_{1,2}=1$ and $E_{1,1}=E_{2,1}=E_{2,2}=0$. We may then view $\mathcal{E}_{(2),B_{s}}$ as factoring as $\mathbb{G}_{a(2),k(\underline{B})}\to\mathbb{G}_{a(2),k(\underline{B})}\subset SL_{2(2),k(\underline{B})}$.
To compute $JT_{sl_{2},2,S(\lambda)}^{exp}(\underline{B})$, we identify the Jordan type of
$(\mathcal{E}_{(2),B_{0}})_{*}(u_{1})+(\mathcal{E}_{(2),B_{1}})_{*}(u_{0})=a_{0}E\cdot u_{1}+a_{1}^{p}E\cdot u_{0}\in k(\underline{B})\otimes kSL_{2(2)}$
on $S(\lambda_{0})\otimes S(\lambda_{1})^{(1)}$. Here, the action of $a_{0}E\cdot u_{1}$ on $S(\lambda_{0})$ is zero, whereas the action of $a_{1}^{p}E\cdot u_{0}$ on $S(\lambda_{0})$ is $a_{1}^{p}$ times the action of $E\in sl_{2}$ on $S(\lambda_{0})$; the action of $a_{0}E\cdot u_{1}$ on $S(\lambda_{1})^{(1)}$ is $a_{0}$ times the action of $E\in sl_{2}$ on $S(\lambda_{1})$ and the action of $a_{1}^{p}E\cdot u_{0}$ on $S(\lambda_{1})^{(1)}$ is zero.
Observe that the Jordan type of $E$ on $S(\lambda_{0})=[m]$ where $m=\lambda_{0}+1$ and on $S(\lambda_{1})=[n]$ where $n=\lambda_{1}+1$. Thus, if $a_{0}\not=0,\ a_{1}=0$, then the Jordan type $S(\lambda)$ is $m\cdot[n]$; and if $a_{0}=0,\ a_{1}\not=0$, then the Jordan type $S(\lambda)$ is $n\cdot[m]$. More generally, $S(\lambda_{0})\otimes S(\lambda_{1})^{(1)}$ as a $k(\underline{B})\otimes k[u]/u^{p}$-module with $1\otimes u$ acting as $a_{0}E\cdot u_{1}+a_{1}^{p}E\cdot u_{0}$ can be identified with a tensor product of two $k[u]/u^{p}$-modules: the first with with $u$ acting as $a_{1}^{p}E\cdot u_{0}$ on $S(\lambda_{0})$, the second with $u$ acting as $a_{0}E\cdot u_{1}$ on $S(\lambda_{1})^{(1)}$. Consequently, if both $a_{0},\ a_{1}$ are non-zero, then the the Jordan type of $a_{1}^{p}E\cdot u_{0}+a_{0}E\cdot u_{1}$ on $S(\lambda)$ equals $m]\otimes[n]$.
@@
Let $V$ be a polynomial representation of $GL_{N}$ of degree <p<p and consider the $i$-th Frobenius twist $V^{(i)}$ for some $i\geq 0$. We fix some $r>i$ and proceed to determine $JT_{\mathfrak{g}l_{N},r,V^{(i)}}(\underline{B})$ for $\underline{B}=(B_{0},\ldots,B_{r-1})$, an $r$-tuple of $p$-nilpotent, pairwise commuting elements of $k(\underline{B})\otimes\mathfrak{g}l_{N}$. Then the action of $(\exp_{B_{s}})_{*}(u_{r-s-1})$ on $V^{(i)}$ is non-zero if and only if i=rs1i=r-s-1, in which case this action equals the action of $B_{s}\in k(\underline{B})\otimes\mathfrak{g}l_{N}$ acting on $k(\underline{B})\otimes V$. Consequently, $JT_{\mathfrak{g}l_{N},r,V^{(i)}}^{exp}(\underline{B})=JT(B_{s},V)$ where s=ri1s=r-i-1.
More generally, we can take a polynomial representation $V_{s}$ of $GL_{N}$ for each $s,\ 0\leq s<r$ with each $V_{s}$ of degree <p<p and consider $M=\otimes_{s=1}^{r-1}V_{s}^{(s)}$. Let $\underline{B}$ be a geometric point of $\mathcal{C}_{r}(\mathfrak{g}l_{N})$. The action of $u=\sum_{s=0}^{r-1}(exp_{B_{s}})_{*}(1\otimes u_{r-s-1})$ on $k(\underline{B})\otimes M$ is the tensor product of the actions of $B_{s}\in k(\underline{B})\otimes\mathfrak{g}l_{N}$ acting on $k(\underline{B})\otimes V_{s}$.
Thus, $JT_{\mathfrak{g}l_{N},r,M}^{exp}(\underline{B})\ =\ \bigotimes_{s=0}^{r-1}JT(B_{r-s-1},V_{s}).$
@@
We employ Theorem 3.7 of [16] to consider the general situation of $(\mathbb{G},\mathcal{E}_{(r)})$ an affine group scheme of exponential type of height $r$ and $M$ a $\mathbb{G}$ module of dimension $m$. Namely, if $\rho:\mathbb{G}\to GL_{m}$ determines $M$, then $(\rho\circ\mathcal{E}_{(r),B_{s}})_{*}(1\otimes u_{r-s-1})$ acting on $k(\underline{B})\otimes M$ is given by the $m\times m$ matrix whose $(i,j)$-entry equals the coefficient of $t^{p^{r-s-1}}\in k(\underline{B})\otimes\mathbb{G}_{a(r)}$ of $(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1})$ applied to $\rho^{*}(X_{i,j}).$
Thus $JT^{exp}_{\mathfrak{g},r,M}(\underline{B})$ is the Jordan type of the sum (indexed by $s,0\leq s<r$) of these $m\times m$ matrices.
6. Vector bundles and stratifications
In this section, we revisit the construction of vector bundles associated to $\mathbb{G}_{(r)}$-modules of constant Jordan type (and more general constant $j$-type). We refer to [7] for an introduction to modules of constant Jordan type and for numerous examples of these interesting $\mathbb{G}$-modules. In [16], such vector bundles were associated to $\mathbb{G}_{(r)}$-modules of constant Jordan type in order to provide invariants which can distinguish some $\mathbb{G}_{(r)}$-modules whose Jordan type functions are equal.
As we point out, the constructions of [16] when applied to an arbitrary $\mathbb{G}_{(r)}$-module $M$ yield coherent sheaves on $\mathbb{P}V_{r}(\mathbb{G})$ whose restrictions to each stratum associated to $JT_{\mathbb{G},r,M}(-)$ are locally free.